# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 8. Limitations of Existing Works and Future Research Direction
Existing attack methods have some limitations. For instance, the DAN attack (shen2023anything, ) builds on jailbreak prompts gathered over six months, extending from the inception of ChatGPT-related sources to May 2023. It is recognized that adversaries have the potential to persist in refining jailbreak prompts for specific objectives beyond the documented collection timeframe. This demonstrates the dynamic nature of adversarial strategies, with the understanding that new, optimized prompts may emerge even after the initial data collection phase. Moreover, most of the existing studies are primarily performed on ChatGPT (dengmasterkey, ; shen2023anything, ; li2023multi, ). It remains unclear whether potential vulnerabilities exist in other LLMs, such as Vicunna (chiang2023vicuna, ), Bard (manyika2023overview, ), and Bing Chat (bingchat, ). For example, MASTERKEY (deng2023jailbreaker, ) may not achieve comparable performance on Vicunna as on ChatGPT. For prompt-based attacks, manual prompts are often more effective than automated prompting-based attack methods. As automated attack methods are primarily designed for generalized tasks, they may not work as effectively for specific tasks as manually crafted prompts. The unlabeled and imbalanced real-world data further complicate the development of effective automated prompting-based attack methods (shin2020autoprompt, ). For backdoor attacks, most of them are focused on classification or similar tasks (cai2022badprompt, ), e.g., sentiment analysis, and opinion classification. More attention should be paid to other NLP tasks, including question answering, text summarization, and language translation.
@@
The underlying philosophy behind privacy attacks lies in the correlation between the level of accessibility an adversary possesses and its ability to extract sensitive information or exert control over target victim LLMs. More access leads to a greater potential for the adversary to recover sensitive data or influence the target LLM (li2023privacy, ). For instance, when the adversary has access only to the black-box model, the attacker might be able to leverage training data extraction attacks to recover a limited set of private data. However, if the adversary is granted white-box information such as gradients, the attacker can leverage this extra information to accurately recover private data samples (deng2021tag, ). This expanded access can facilitate various privacy attacks, including attribute inference, embedding inversion, and gradient leakage attacks. In gradient-based attacks, the adversary needs the white-box information of a model, which is sometimes impractical. Moreover, most of the existing privacy attacks are designed for vision models. A limited number of studies have reported gradient leakage attacks specifically on language models, e.g., the LAMP attack (balunovic2022lamp, ). In essence, the increased access empowers adversaries to perform more sophisticated and targeted privacy attacks against LLMs, potentially compromising sensitive information or gaining access to the internal model architecture. The early MIAs were based on the white-box access assumption (shokri2017membership, ), which is sometimes impractical in real-world deployment. The evaluation dataset for some attacks, e.g., ProPILE was built solely from private information available in open-source datasets provided by major corporations, ensuring ethical data acquisition (kim2023propile, ). However, it is crucial to note that the heuristic data collection process might potentially lead to instances of bias, disassociation, or noise. This adds uncertainty and potential inaccuracies in the benchmark dataset, requiring attention when interpreting the results.
@@
However, most attack methods (e.g., BadGPT, BadPrompt, and Trojaning attacks) described in the existing studies are designed for only relatively small NLP models. Only a few are tested on LLMs (e.g., ProPILE, DAN, and JAILBREAKER). Also, the high cost of accessing commercialized LLMs, such as GPT-3.5 or upper versions, contributes to the lack of attack evaluations on LLMs. Besides, the in-depth vulnerability analysis in terms of privacy attacks and security issues is still lacking for LLMs. One of the reasons can be attributed to the limited number of performance evaluation metrics (e.g., perplexity) in the language domain to comprehensively evaluate attack and defense effectiveness. Also, in the FL environment, it requires very high computational power to train LLMs with such large datasets (sutitanic, ). It is an open research challenge to develop effective attack and robust defense methods along with the proper evaluation techniques for LLMs.
@@
For defense, studies reported that ChatGPT’s safety protections are good enough to prevent single jailbreaking prompts but still vulnerable to multi-step jailbreaking (li2023multi, ). Moreover, the new Bing AI chatbot (bing, ) is more vulnerable to these direct prompts. System-mode self-reminder defense techniques are inspired by the human-like reasoning capabilities of LLMs (xi2023defending, ). The more discerning question regarding LLM reasoning processes with or without self-reminder remains unsolved. To acquire a comprehensive understanding of the reasoning processes of large neural networks, more in-depth investigation is essential. Although the side effects of self-reminder have been explored on typical user queries across various NLP tasks, evaluating its effect on any type of user query poses a challenge, making it difficult to fully understand its impact on user experience (wu2023defending, ). Considering the shortcomings mentioned, developing more flexible self-reminding systems and expert frameworks that improve safety, trustworthiness, and accountability in LLMs without compromising effectiveness can be a fundamental research challenge to protect LLMs from jailbreaking attacks. Moreover, individuals with malicious intent are highly active in online forums, sharing and discussing new strategies. Frequently, they keep these exchanges private to evade detection. Consequently, it is essential to conduct further research and studies aimed at identifying and implementing effective defense strategies to mitigate the risks posed by the latest jailbreaking attacks. Efficient strategies for defending against backdoor attacks in a black box environment are still lacking (kandpal2023backdoor, ). Existing defense mechanisms (shao2021bddr, ), for specific learning tasks in LMs are not evaluated for the other learning tasks like, text summarizing, and prompt-based learning. Moreover, it is found in the literature that prompt-based PLMs are highly susceptible to textual backdoor attacks (xu2022exploring, ), (du2022ppt, ). Addressing the challenge of textual backdoor attacks in prompt-based paradigms, particularly in the few-shot learning setting (wang2020generalizing, ), is another unresolved challenge. MDP (Masking-Differential Prompting) defense (xi2023defending, ) faces challenges in various NLP tasks like paraphrasing and sentence similarity (gao2020making, ). While MDP has demonstrated strength against earlier backdoor attacks, it may not be effective against more recent attacks like BadPrompt (cai2022badprompt, ) and BToP (xu2022exploring, ). Perplexity-based methods and filtering-based methods may not work well when attackers use synonymous trigger keys (huang2023composite, ). Moreover, developing dynamic defense methods considering the above factors is a challenging future task. Currently, the predominant focus of investigation on backdoor attacks revolves around text classification in LLMs. However, a notable gap exists in the literature concerning investigations into backdoor attacks on various tasks for which LLMs find widespread application, e.g., text summarization and text generation (yang2023comprehensive, ). Understanding and addressing backdoor attacks in various tasks for which LLMs are employed is crucial for developing effective defense mechanisms and ensuring secure deployment of LLMs. While poisoning attacks on ML models have been investigated in the literature (mooreFL2023, ), there is not yet an effective solution for several attack methods, including ProAttack(zhao2023prompt, ) and Badprompt(cai2022badprompt, ). Further research in diverse tasks and models can enhance the knowledge and understanding of the security impacts of LLMs, as well as facilitate the development of robust and trustworthy LLM systems. Defense techniques, such as dataset cleaning, and removing near duplicate poisoned samples and anomalies sometimes slow down the model development process in order to defend against data poisoning attacks. Other defense methods, e.g., stopping training after certain epochs achieve a moderate defense against poisoning attacks but degrade the model utility (wan2023poisoning, ).
Gradient perturbation (hu2023defenses, ) and DP-SGD-based methods (abadi2016deep, ) are frequently used to defend against privacy attacks in LLMs. It can prevent the private training data from being leaked based on the parameter configurations at a small cost of model utility. Limiting the accessibility to the model and generating limited prediction results might be another option (zanella2020analyzing, ). Extensive research studies can obtain proper knowledge of to what extent algorithmic defenses such as differential privacy can prevent PII disclosure without compromising model utility. In the post-processing phase, for API-access models such as GPT-3, it is advisable to integrate a detection module that examines the output text to identify sensitive information. If sensitive content is detected, the system should either decline to provide an answer or apply masks to safeguard the sensitive information (huang2022large, ). Also, for image models, a recent study has demonstrated that adding a standard level of random noise into the gradient update might not always work well to prevent gradient leakage attacks on medical images (das2023privacy, ).
Considering the above limitations of existing defense techniques in LLMs, developing a defense mechanism for these privacy attacks for LLMs would be an imperative task. Secure multi-party computation (cramer2015secure, ) can be another way to defend against privacy attacks in LLMs, which can be explored in future research. An ideal defense method should be able to effectively achieve a balance between model utility and privacy protection, incorporate real-time privacy monitoring, and demonstrate resilience against evolving attacks, such as BadPrompt, multi-step jailbreaking, and ProAttack. This necessitates the exploration of multi-modal attacks, robust detection, and prevention-based techniques to comprehensively understand privacy risks and protect privacy in evolving threat landscapes.

### v2

@@ 8. Limitations of Existing Works and Future Research Direction
Existing Attack Methods, their Limitations, and Future Research Direction: In LLMs, various categories of security and privacy attacks have emerged, posing significant risks to LLM systems. Among security attacks, prompt injection is a prominent technique in which attackers craft malicious prompts, either manually (Liu et al., 2023a) or automatically (Shin et al., 2020), to mislead the LLM into generating inappropriate or harmful outputs. These prompts are designed to bypass the model’s safety alignments (Jiang et al., 2023b), enabling the generation of content as per the attacker’s intent. Jailbreaking attacks involve creating malicious prompts in tricky way (Spennemann, 2023), such as character role-play (Gupta et al., 2023) and attention shifting (Liu et al., 2023b), so that the LLM generates inappropriate/harmful contents. Data poisoning attacks involve the insertion of malicious data samples during the model’s training/fine-tuning phases (Wan et al., 2023; Kurita et al., 2020), introducing biases or vulnerabilities that compromise the model’s functionality. Similarly, backdoor attacks introduce hidden “backdoors” (via trojaning) during the training/fine-tuning phase (Shi et al., 2023b), which are activated by the presence of backdoor triggering words in the prompts, making the LLM system vulnerable. LLM privacy attacks focus on extracting sensitive information such as private training/fine-tuning data, personal information, and even model components (e.g., gradients or model architecture). Techniques used in these attacks include gradient leakage attacks (Deng et al., 2021), MIA (Shokri et al., 2017), and PII leakage attacks (Nakka et al., 2024).
@@
Recent studies have shown that existing attack methods have some drawbacks. For instance, the DAN attack (Shen et al., 2023) builds on jailbreak prompts gathered over six months, extending from the inception of ChatGPT-related sources to May 2023. It is recognized that adversaries have the potential to persist in refining jailbreak prompts for specific objectives beyond the documented collection time-frame. This demonstrates the dynamic nature of adversarial strategies, with the understanding that new, optimized prompts may emerge even after the initial data collection phase. Moreover, most of the existing studies on jailbreaking attacks are primarily focused on ChatGPT (Deng et al., 2023; Shen et al., 2023; Li et al., 2023b). It remains unclear whether potential vulnerabilities exist in other LLMs, such as Vicuna (Chiang et al., 2023), Bard (Manyika and Hsiao, 2023), and Bing Chat (Mehdi, 2023). For example, MasterKey (Deng et al., 2023) may not achieve comparable performance on Vicuna as it does on ChatGPT. For prompt-based attacks, manual prompts are often more effective than automated prompt-based attack methods. As automated attack methods are primarily designed for generalized tasks, they may not demonstrate same effectiveness for specific tasks as manually crafted prompts. The unlabeled and imbalanced real-world data further complicate the development of effective automated prompting-based attack methods (Shin et al., 2020). For backdoor attacks, most of them are focused on classification or similar tasks (Cai et al., 2022), e.g., sentiment analysis and opinion classification. More attention should be paid to the attacks that perform other NLP tasks, including question answering, text summarization, and language translation.
@@
The underlying philosophy behind privacy attacks lies in the correlation between the level of accessibility an adversary holds and its ability to extract sensitive information or exert control over target victim LLMs. More access leads to a broader potential for the adversary to recover sensitive data or influence the target LLM (Li et al., 2023a). For instance, when the adversary has access only to the black-box model, the attacker might be able to leverage training data extraction attacks to recover a limited set of private data. However, if the adversary is granted white-box information such as gradients, the attacker can leverage this extra information to accurately recover more private training instances (Deng et al., 2021). This expanded access can facilitate various privacy attacks, including attribute inference attacks, embedding inversion, and gradient leakage attacks. In gradient-based attacks, the adversary needs the white-box information of a model, which is sometimes impractical in real-world practice. Moreover, most of the existing privacy attacks are designed for vision models. A limited number of studies have reported gradient leakage attacks specifically on language models, e.g., the LAMP attack (Balunovic et al., 2022). In essence, the increased access empowers adversaries to perform more sophisticated and targeted privacy attacks against LLMs, potentially compromising sensitive information or gaining access to the internal model architecture. The early MIAs were based on the white-box access assumption (Shokri et al., 2017), which is sometimes impractical in real-world deployment. The evaluation dataset for some attacks, e.g., ProPILE was built solely from private information available in open-source datasets provided by major corporations, ensuring ethical data acquisition (Kim et al., 2024). However, it is crucial to note that the heuristic data collection process might potentially lead to instances of bias, disassociation, or noise. This adds uncertainty and potential inaccuracies in the benchmark dataset, requiring attention when interpreting the results.
@@
However, most attack methods (e.g., BadGPT, BadPrompt, and Trojaning attacks) described in the existing studies are designed for only relatively small NLP models. Only a few are tested on LLMs (e.g., ProPILE, DAN, and JAILBREAKER). Also, the high cost of accessing commercialized LLMs, such as GPT-3.5 or upper versions, contributes to the lack of attack evaluations on LLMs. Besides, the in-depth vulnerability analysis in terms of privacy attacks and security issues is still lacking for LLMs. One of the reasons can be attributed to the limited number of performance evaluation metrics (e.g., perplexity) in the language domain to comprehensively evaluate attack and defense effectiveness. Also, in the FL environment, it requires very high computational power to train LLMs with such large datasets (Su et al., 2023). It is an open research challenge to develop effective attack and robust defense methods along with the proper evaluation techniques for LLMs. The correctness of LLM-generated content has always been a major concern in this area of research. Due to the knowledge gaps, i.e., missing or outdated information might always be present in LLMs. The problem of hallucination in LLM has been investigated and evaluated from the knowledge-gap perspective but is yet to be investigated from other perspectives such as safety, i.e., abstaining to generate harmful content (Chang, 2023).
For LLMs, comprehensive exploration of their vulnerabilities under security and privacy attacks remains an essential area of study. Future research should examine the applicability of well-established attack methods from other domains, such as computer vision, to LLMs. In addition to existing methods, novel attack techniques should be developed to comprehensively inspect various vulnerabilities of LLMs, potentially targeting multiple objectives with a single attack approach. Moreover, special attention should be given to developing appropriate metrics for evaluating the impact of vulnerabilities related to security and privacy attacks in LLMs. Explainable AI (XAI) can play a vital role in this domain by increasing transparency and explainability within attack systems, allowing for a better understanding of LLM vulnerabilities. Developing XAI techniques to enhance the interpretability of LLM vulnerabilities is another essential research direction for advancing the security and privacy of LLMs.
Existing Defense Mechanisms, Challenges, and Future Research Direction: To mitigate vulnerabilities posed by security and privacy attacks, various defense techniques have been proposed in existing research. Instruction defense (Schulhoff, 2024a), paraphrasing (Jain et al., 2023), and re-tokenization (Jain et al., 2023) methods are commonly employed to defend against prompt injection attacks. To defense jailbreaking attacks, SmoothLLM (Robey et al., 2023), LLM Guard (AI, 2024), and perplexity filtering (Gonen et al., 2022) are widely used. Data curation techniques (Yan et al., 2024a), including filtering poisoned content (Wan et al., 2023) and detecting trigger-embedded inputs (Chen et al., 2018), are widely used to mitigate data poisoning attacks and backdoor attacks. On the other hand, DP-SGD (Abadi et al., 2016) is a widely adopted technique to defend against gradient leakage attacks, MIAs, and PII leakage attacks. In particular, noise perturbation (Zhu et al., 2019) is a common strategy to defend against gradient leakage attacks. Various techniques are prevalent to mitigate MIA in LLMs, including model pruning (Liu et al., 2018a) and knowledge distillation (Hinton, 2015). To prevent PII leakage attacks, training data curation (Yan et al., 2024a) and restricting attacker access to training data or model (Inan et al., 2021b) serve as viable countermeasures.
For defense, studies reported that ChatGPT’s safety protections are good enough to prevent single jailbreaking prompts however, it is still vulnerable to multi-step jailbreaking (Li et al., 2023b). Moreover, the new Bing AI chatbot (Microsoft, 2023) is more vulnerable to these direct prompts. System-mode self-reminder defense techniques are inspired by the human-like reasoning capabilities of LLMs (Xi et al., 2023b). The more discerning question regarding LLM reasoning processes with or without self-reminder remains unsolved. To acquire a comprehensive understanding of the reasoning processes of large neural networks, more in-depth investigation is highly required. Although the side effects of self-reminder have been explored on typical user queries across various NLP tasks, evaluating its effect on any type of user query poses a challenge, which makes it difficult to fully understand its impact on user experience (Wu et al., 2023c).
@@
Considering the shortcomings mentioned, developing more flexible self-reminding systems and expert frameworks that improve safety, trustworthiness, and accountability in LLMs without compromising effectiveness can be a fundamental research challenge to protect LLMs from jailbreaking attacks. Furthermore, individuals with malicious intent are highly active in online forums, sharing and discussing new strategies. Frequently, they keep these exchanges private to evade detection. Consequently, it is essential to conduct further research and studies aims to identify and implement effective defense strategies to mitigate the risks posed by the latest jailbreaking attacks. Efficient strategies for defending against backdoor attacks in a black-box environment are still lacking (Kandpal et al., 2023). Existing defense mechanisms ((Sur et al., 2023), (Qi et al., 2020), (Shao et al., 2021)), for specific learning tasks in LMs are not evaluated for the other learning tasks like, text summarizing, and prompt-based learning. Moreover, it is found in the literature that prompt-based PLMs are highly susceptible to textual backdoor attacks (Xu et al., 2022), (Du et al., 2022). Addressing the challenge of textual backdoor attacks in prompt-based paradigms, particularly in the few-shot learning setting (Wang et al., 2020), is another unresolved challenge. MDP (Masking-Differential Prompting) defense (Xi et al., 2023b) faces challenges in various NLP tasks like paraphrasing and sentence similarity (Gao et al., 2020). Its performance under few-shot learning remains uncertain due to a lack of practical evaluation. While MDP has demonstrated strength against earlier backdoor attacks, it may not be effective against recently introduced attacks like BadPrompt (Cai et al., 2022) and BToP (Xu et al., 2022). Perplexity-based methods and filtering-based methods may not work well when attackers use synonymous trigger keys (Huang et al., 2023b). Furthermore, developing dynamic defense methods considering the above factors is a challenging future task. Currently, the predominant focus of investigation on backdoor attacks revolves around text classification in LLMs. However, a notable gap exists in the literature concerning investigations into backdoor attacks on various tasks for which LLMs find widespread application, e.g., text summarization and text generation (Yang et al., 2023). Understanding and addressing backdoor attacks in various tasks for which LLMs are employed is crucial for developing effective defense mechanisms and ensuring secure deployment of LLMs. While poisoning attacks on ML models have been investigated in the literature (Moore et al., 2023), there is not yet an effective solution for several attack methods, including ProAttack(Zhao et al., 2023a) and Badprompt(Cai et al., 2022). Further research in diverse tasks and models can enhance the knowledge and understanding of the security impacts of LLMs, as well as facilitate the development of robust and trustworthy LLM systems. Defense techniques, such as dataset cleaning, and removing near duplicate poisoned samples and anomalies, sometimes slow down the model development process in order to defend against data poisoning attacks. Other defense methods, e.g., stopping training after certain epochs, achieve a moderate defense against poisoning attacks but degrade the model utility (Wan et al., 2023). Gradient perturbation (Hu et al., 2023b) and DP-SGD-based methods (Abadi et al., 2016) are frequently used to defend against privacy attacks in LLMs. It can prevent the private training data from being leaked based on the parameter configurations at a small cost of model utility. Limiting the accessibility to the model and generating limited prediction results might be another option (Zanella-Béguelin et al., 2020). Extensive research studies can obtain proper knowledge of to what extent algorithmic defenses such as differential privacy can prevent PII disclosure without compromising model utility. In the post-processing phase, for API-access models such as GPT-3, it is advisable to integrate a detection module that examines the output text to identify sensitive information. If sensitive content is detected, the system should either decline to provide an answer or apply masks to safeguard the sensitive information (Huang et al., 2022). Also, for image models, a recent study has demonstrated that adding a standard level of random noise into the gradient update might not always work well to prevent gradient leakage attacks on medical images (Das et al., 2023). Recently, an open language model (OLMo) has been introduced to provide open access to the data, code, and model (Soldaini et al., 2023). The main purpose of OLMo is to facilitate open research on language models. They performed PII filtering to remove it from the data, and they also provided a tool to remove PII data upon request. Though it followed existing practices to obscure the PII exposure and identify and remove toxic content, it did not explicitly discuss the impacts of various attacks outlined in this survey paper and how to mitigate those risks. Secure multi-party computation (Cramer et al., 2015) can be another way to defend against privacy attacks in LLMs, which can be explored in future research endeavors. Considering the above limitations of existing defense techniques in LLMs, developing a defense mechanism for these privacy attacks for LLMs would be an imperative task.
Future research initiatives for enhancing security and privacy in LLMs can be directed toward several key areas. An ideal defense method should be able to effectively achieve a balance between model utility and security & privacy protection. The development of real-time privacy monitoring systems is essential to improve the resilience of privacy-preserving LLMs. This necessitates the exploration of robust detection techniques against various security and privacy attacks. Furthermore, a thorough evaluation of less-explored defense techniques, such as secure multi-party computation (SMPC), is necessary to assess their effectiveness against LLM vulnerabilities. Finally, leveraging the capabilities of XAI can improve transparency in LLM defense mechanisms.
Role of explainable AI in Enhancing Security and Privacy of LLMs: LLM’s ability to make decisions on various learning tasks is often criticized due to its black-box nature, e.g., non-interpretable weights. It makes it more challenging for the new practitioners and developers of this field, as it hinders their ability to clearly interpret and understand its application, particularly in critical cases. Explainable AI (XAI) can help to bridge this gap by developing methods to interpret and explain complex LLM systems, the decision-making process, and outputs (Cambria et al., 2024). Conducting studies and research to build explainable methods is highly essential to make the usage of LLMs more reliable and trustworthy. XAI can provide transparent insights into the inference process and the dynamic weight assignment by the attention mechanism of the LLMs, which enhances interpretability by highlighting the most influential input features that contribute to its predictions. This explainability makes the model more trustworthy and ensures that its output is explainable. Additionally, it facilitates error analysis and bias detection. XAI has shown its potential in several real-world AI applications such as anomaly detection (30dascoding, 2024; Ali, 2024), cyber-security (Saarela and Podgorelec, 2024), and enhancing data privacy (Ezzeddine, 2024). In ML, several popular methods exist such as Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016), and SHapley Additive exPlanations (SHAP) (Lundberg, 2017), which offer insights into model behavior without requiring access to internal model components. These methods analyze the association between inputs and outputs to identify features that most significantly contribute to model predictions. The explainability of code generation models such as CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020) were proposed to understand code syntax and semantics (Yang et al., 2024a). Also, the ad-hoc explanation further clarifies the model’s decision (Lipton, 2018).
The role of XAI in the security and privacy aspect of LLM is crucial. Specifically in its development phase, having prior knowledge about the attackers’ ability (e.g., black-box access to the model, injecting poisonous instances to the training/fine-tuning data) and the vulnerable components (training/fine-tuning data and the model itself) of LLM systems will assist in developing robust techniques against security and privacy attacks. However, XAI may expose critical aspects of the LLMs, including their architecture, components, and the dynamic weight allocation by attention mechanism that contribute to predictions. Such disclosures may increase the LLM’s vulnerability to security and privacy attacks. For example, revealing the white-box nature of the model may facilitate the attacker’s access to the model, increasing the risk of gradient leakage attacks or MIA (Ezzeddine, 2024). Recent studies have claimed that explainability in LLMs may raise further security concerns, especially with insidious backdoor attacks (Cheng et al., 2023). Lin et al. proposed an XAI approach that can identify the triggers (backdoor attack) that mislead the model to contribute to error classification (Lin et al., 2021). Li et al. developed an XAI technique to identify and quantify the influence of raw data features on successful MIAs. This approach analyzes the data distribution to identify the influential neurons contributing to compromising private data and subsequently trains an MIA ensemble model using attack features derived from the selected neurons (Lin et al., 2021). XAI techniques in image models (e.g., Grad-CAM (Selvaraju et al., 2017)) may compromise model privacy under various attacks, such as gradient leakage attacks. In some cases, XAI enables more accurate reconstruction of private training data compared to models that rely solely on predictions (Zhao et al., 2021). XAI-aware model extraction attack (XaMEA) was proposed to exploit spatial knowledge from decision explanations (Yan et al., 2023). It illustrated that the transparency provided by XAI may facilitate the attacker’s access to the model, making it easier to exploit it. This increased explainability can make the model more vulnerable to model extraction attacks compared to prediction-only models. However, the vulnerabilities mentioned above are mostly explored for DNNs. In LLMs, the vulnerabilities in the XAI context have been explored to a limited extent. Moreover, there are several unique characteristics of LLMs that are different from DNNs, e.g., large-scale data and model parameters, task-agnostic, and semantic language understanding, which makes it more challenging to design XAI methods for better interpretations. The comprehensive interpretation of LLM vulnerabilities under XAI, including backdoor attacks, membership inference attacks (MIA), and model extraction attacks, has yet to be fully explored. Additionally, vulnerabilities related to prompt injection and jailbreaking attacks in the context of XAI still remain unexplored. Furthermore, open-sourced LLMs are publicly accessible and may provide greater explainability than closed-source models. However, this enhanced accessibility can make them more vulnerable to attacks when XAI techniques are employed. Thus, these LLMs may be more susceptible to attacks due to XAI. Further research is essential to fully understand the extent and impact of attackers’ capabilities under XAI. Additionally, mitigation techniques should be developed in accordance with the identified risks and impacts.